Find all points on y-axis which are 5 units from point (-5,3)
Answer by stanbon(75887)

Find all points on y-axis which are 5 units from point (-5,3)
-----
All points on the y-axis look like (0,y)
-------
Equation:
(0,y) to (-5,3) = 5 units
--------------------------------
sqrt[(y-3)^2 + (0+5)^2)] = 5
-----
y^2-6y+9 + 25 = 25
----------------------
y^2-6y +9 = 0
-----
(y-3)(y-3) = 0
y = 3
------
------------------------
Ans: (0,3)
------------------------
Cheers,
Stan H.
-------------